# Binary Stars Approaching Supermassive Black Holes: Hydrodynamics of Stellar Collisions, Mass Fallback and Partial TDEs

### v1

@@ 3 Results
We now examine in detail the properties of the remnants for the four representative simulation runs, along with the orbital characteristics of the mass-loss debris and the fallback rate onto the central SMBH.
@@ 3.1 Merger Remnant(s)
Figure 5: Mass loss fraction as a function of time for Runs 1 & 3. The lines with the $\bullet$” markers of different colors represent results with different particle numbers ($N=5,1,10\times 10^{5}$ per star, Runs 1a, 1b, 1c). The line with “$\times$” markers represents the result for Run 3. Note that $t=0$ roughly indicates the moment of collision. $M_{\rm tot}$ represents the total mass of the initial binary.
In Section 2.2, we describe how we identify the SPH particles that belong to the merger remnant(s). Figure 5 presents the results of precisely calculating the mass loss fraction for two head-on collisions using the Bernoulli method. It is evident that the fates of most particles - whether they remain part of the merger remnant or are lost - are determined quickly after the collision. This occurs more rapidly in collisions during gentle encounters compared to those in deep encounters, as the higher impact velocity in deep encounters results in a more violent interaction among fluid elements, leading to a slower mass loss process.
@@
For head-on collisions in both deep and gentle encounters, the mass loss fraction stabilizes at approximately 5%.
Figure 6: The density profile of the merger remnant (left column) and the fraction of the total mass enclosed within a given radius relative to the total mass of the merger remnant (right column) are shown for Run 1a (upper row) and Run 3 (lower row). The solid lines represent the profiles at different times after the collision, while the dashed line corresponds to a $2M_{\odot}$ $\gamma=5/3$ polytrope with radius determined by the mass-radius relation $R_{\star}\propto M_{\star}^{0.8}$, which also approximates the structure of a main-sequence star. The insets in the right panels highlight the mass distribution at the outskirts of the remnant.
In addition to the mass loss, we are also interested in the structural properties of the merger remnant. Clearly, not all particles settle down immediately after the collision, and some asymmetry exists within the merger remnant during the short post-collision phase. For now, we ignore the effects of this asymmetry. Starting from the remnant’s center of mass, we calculate the density $\rho_{\rm rem}$ at different radius $r_{\rm rem}$ by dividing the mass and volume of spherical shells. We also compute the fraction of the total mass enclosed within a given radius $M_{\rm rem}(<r_{\rm rem})$ relative to the total mass of the merger remnant, as determined using the Bernoulli method. Figure 6 illustrates the structure of the merger remnant following the head-on collision (Runs 1 & 3).
@@
Compared to a typical $2M_{\odot}$ main-sequence star (modeled here as a $\gamma=5/3$ polytrope, with its radius determined by the mass-radius relation $R_{\star}\propto M_{\star}^{0.8}$; represented by the dashed line in Figure 6), the merger remnant exhibits a more concentrated core and an extremely extended envelope. Over 5% of the total mass is distributed in the envelope at $r_{\rm rem}>5R_{\odot}$.
We can also calculate the total energy of the remnant based on the SPH simulation output. Taking the head-on collision in the deep encounter (Run 1a) as an example, we compute the kinetic energy $E_{k,\rm rem}$ relative to the center of mass, gravitational potential energy $E_{p,\rm rem}$, and internal energy $U_{\rm rem}$ for all SPH particles belonging to the remnant. These components were then used to determine the total energy $E_{tot,\rm rem}$ of the remnant. The results are presented in Figure 7.
Figure 7: Energy of the merger remnant in Run 1a as a function of time. The different solid lines represent the kinetic energy $E_{k,\rm rem}$ (relative to the remnant center of mass), gravitational potential energy $E_{p,\rm rem}$, internal energy $U_{\rm rem}$, and total energy $E_{\rm tot,\rm rem}$.
@@
For the grazing collisions (Runs 2 & 4), the larger impact parameters typically prevent the two stars from merging during the first collision. In the case of the deep encounter (Run 2), the tidal forces from the SMBH have already disrupted the binary energy-wise. After the first collision, the two stars do not collide again but instead undergo a process resembling the standard binary tidal breakup: one star becomes bound to the SMBH’s orbit, while the other is ejected. However, both stars have undergone significant perturbations, leaving their structures vastly different from those of typical main-sequence stars.
@@
Figure 9: Same as Figure 6, but for Runs 2 and 4. Note that the “bump” in the blue line in the upper left panel represents the other remnant. When two remnants are present, the exact remnant mass cannot be easily determined. Therefore, we calculate the fraction of the total mass enclosed within a given radius relative to the initial total mass of the system (which is $1M_{\odot}$ for Run 2 and $2M_{\odot}$ for Run 4).
We also analyze the property of the merger remnant following a grazing collision (Runs 2 & 4), with the results shown in Figure 9. Unlike the cases of head-on collisions, for the deep encounter, we select the remnant that is ultimately bound to the SMBH and compared it to a $1M_{\odot}$ $\gamma=5/3$ polytrope. Additionally, when calculating the fraction of mass enclosed within different radii, we no longer use the merger remnant’s total mass determined via the Bernoulli method. Instead, we directly adopt the initial mass ($2M_{\odot}$ for the gentle encounter and $1M_{\odot}$ for the deep encounter). We also include a red dashed line indicating the estimated total mass of the remnant, inferred from the total mass fraction. This suggests that the mass loss fraction is approximately 4% in Run 2 and around 1% in Run 4. The latter is roughly in line with the value obtained using Bernoulli method (see Table 1).
@@ 3.2 Mass Loss Debris and Fallback
In Paper I, we suggested the mass loss from stellar collisions induced by binary-SMBH encounters is mass loss will accrete onto the SMBH, generating accretion luminosity or flares even without a stellar TDE. As shown in Figures 1-5, it is indeed the case that, since the collision velocity is typically comparable to the stellar surface escape velocity (see Figure 9 of Paper I), collisions generally result in a mass loss of approximately 5%. This is consistent with the results of stellar collisions in the absence of an SMBH (e.g., Benz & Hills, 1987; Lai et al., 1993; Freitag & Benz, 2005). A key difference between this mass-loss debris and the typical TDE debris is that the former is not produced by tidal forces, and thus does not remain in the same plane or have the very small launch angles characteristic of the TDE debris (which results in an extremely narrow stream). Instead, the mass-loss debris generated by stellar collisions typically has much larger launch angles, and therefore, in terms of morphology, these debris are more like a “cloud” than a “stream”. Similar to TDEs, because our binary is initially on a parabolic orbit relative to the SMBH, only about half of the mass-loss debris will fall back onto the SMBH.
Since we can identify which SPH particles belong to the mass loss in the case of a head-on collision (Runs 1 & 3), here we will focus our analysis on the head-on collisions in both deep and gentle encounters. For grazing collisions, while the remnant properties differ, the mass-loss debris generally leaves the remnant’s Hill sphere quickly and is dominated by the SMBH’s gravitational influence. Therefore, when considering the fallback onto the SMBH, the results for grazing collisions are similar to those for head-on collisions.
Figure 10: Probability density function (PDF) of the orbital energy of the mass-loss debris relative to the SMBH after head-on collisions in Runs 1a and 3. The results are scaled by $\varepsilon_{\rm col}$, as defined in Eq. (13).
@@
From the Figure 10, we see that the debris energy distribution exhibits a bimodal structure, with a relatively long tail. The bimodal feature arises because the orbital energy of the merger remnant relative to the SMBH is nearly zero, so most particles with nearly zero orbital energy remain bound to the merger remnant. The long tail indicates that many debris particles have relatively high launch velocities relative to the merger remnant. This is due to the shock waves fand fluid forces present during the collision, allowing some debris to achieve higher energies relative to the merger remnant. We also see that in our representative examples, the head-on collision in the deep encounter (Run 1a) results in a wider energy distribution compared to the head-on collision in the gentle encounter (Run 3).
@@
Figure 11: Fallback accretion rate, $\dot{M}(t)$, onto the SMBH from Runs 1 and 3. The different colored lines represent the results for different particle numbers ($N=5,1\times 10^{5}$ per star, Run 1a and 1b). To convert to physical units, here we set $M_{\rm BH}=10^{6}M_{\odot}$, $M_{\star}=1M_{\odot}$ and $R_{\star}=1R_{\odot}$. The black dashed line shows the standard mass fallback rate for a $1M_{\odot}$ star undergoing a TDE, while the red dashed line represents a power-law curve that matches the late-time fallback rate. The dotted line indicates the Eddington limit for a $10^{6}M_{\odot}$ black hole, assuming an accretion efficiency of $\epsilon=0.1$. Here, $t=0$ corresponds to the moment of stellar collision.
@@
The fallback rate curves of the mass-loss debris differ significantly from the standard TDE curve. These curves exhibit a slow rise over an extended period, remain relatively steady, and then follow a power-law decline similar to that of a TDE. The relationship between the curve shape and $\beta_{b}$ is not very clear. The peak fallback rate is much lower than that of a TDE, as the total fallback mass is only a few percent of the fallback mass in a TDE for a star of the same mass. However, the minimum fallback time for the debris is significantly shorter compared to the TDE case, particularly for deep encounters. In gentle encounters, this effect is less pronounced, as the collision occurs long after the pericenter passage. Despite comparable total mass loss, there is a significant difference in the magnitude of the peak fallback rates. In the deep encounter (Run 1), the collision occurs very close to the pericenter, causing the debris to fall back onto the SMBH more quickly. This results in a substantially higher peak rate, which remains above the Eddington limit for several hundred days. In contrast, for the gentle encounter (Run 3), the fallback occurs much more slowly, and the curve consistently stays below the Eddington limit.
@@
We can estimate the approximate size of the final disk by calculating the pericenter distances of the mass-loss SPH particles, based on their angular momenta (relative to the SMBH) after the collision. The results are shown in Figure 12.
@@ 4.1 Summary of Key Results
In this paper, we have conducted a detailed study of the stellar collisions induced by the close encounter between a stellar binary and a SMBH. The binary, with a total mass $m_{12}$, an initial circular orbit, and a semi-major axis $a_{b}$, approaches the SMBH (mass $M_{\rm BH}$) on a parabolic orbit, with a pericenter distance $r_{p}$ from the SMBH. Using SPH simulations, we characterize key properties of stellar collisions and their potential observational consequences.
@@
    Mass loss and remnant(s). Since the collision velocity is comparable to the stellar escape velocity (see Table 1 and Figure 9 in Paper I), some mass loss is expected during the collision. Moreover, depending on the impact parameter (e.g., head-on or grazing collisions), the outcome of the collision may vary. In the case of head-on collisions, the mass loss fraction is approximately $5\%$ (Figure 5), and such collisions consistently produce a single merger remnant. In contrast, grazing collisions exhibit more diverse outcomes. For gentle encounters, while the first collision may not immediately result in a single merger remnant, the binary remains gravitationally bound; the perturbed stars undergo multiple subsequent collisions (Figure 8), ultimately forming a single remnant with minimal mass loss ($\lesssim 1\%$). However, in deep encounters, where the binary system is already disrupted, if the first collision fails to produce a merger, two remnants will form, and the mass loss fraction is relatively higher (around $4\%$).
@@
    Density structure of merger remnant(s). The remnant star formed after a collision exhibits significant differences compared to the original main-sequence star (modeled here as a $\gamma=5/3$ polytrope). The merger remnant features a much higher central density and a highly extended envelope, with approximately $5\%$ of its mass residing beyond five times the original radius (Figs. 6, 9). This extended structure makes the merger remnant more susceptible to partial tidal disruption when it returns to the pericenter around the SMBH. If the two stars do not merge but instead undergo significant dynamical perturbations, the outcomes differ substantially (upper row of Figure 9). These perturbed stars not only develop an extended envelope but also experience a substantial reduction in the central density. Analogous to the Hills mechanism, one of the stars often remains bound to the SMBH. In this case, this perturbed star is also significantly more susceptible to tidal disruption.
@@
    Orbital property of mass-loss debris and mass fallback rate. Nearly half of the mass-loss debris produced during the collision is expected to fall onto the SMBH, undergoing a process analogous to a TDE. From the simulation snapshots (Figs. 1-4), we observe that the “debris cloud” generated by the collision exhibits a morphology distinctly different from the “debris stream” typical of TDEs. Furthermore, an analysis of the binding energy distribution of the debris (Figure 10) reveals significant deviations from the top-hat distribution commonly observed in stellar TDEs. Since the collision does not necessarily occur at the pericenter, unlike in the standard TDE, the resulting mass fallback accretion rate also differs (Figure 11). Although the total mass loss represents only a small fraction of the stellar mass and thus results in a relatively low magnitude of fallback accretion, the shape of the fallback rate curve is notably distinct from that of TDEs. Specifically, it remains approximately constant for an extended period ($\gtrsim 100$ days) before starting to decline (for $M_{\rm BH}=10^{6}M_{\odot}$ and solar-type stars). Additionally, the angular momentum distribution of the mass-loss debris (represented by the debris’ pericenter distance $r_{p}$, see Figure 12) suggests that the resulting accretion disk extend to the original pericenter distance.

### v2

@@ 3 Results: $\gamma=5/3$ polytrope
We now examine in detail the properties of the remnants for the four representative simulation runs for polytropic stars (with $\gamma=5/3$), along with the orbital characteristics of the mass-loss debris and the fallback rate onto the central SMBH.
@@ 3.1 Merger Remnant(s)
Figure 5: Mass loss fraction as a function of time for Runs 1-4. The lines with the $\bullet$” markers of different colors represent results with different particle numbers ($N=5,1,10\times 10^{5}$ per star, Runs 1a, 1b, 1c). Note that $t=0$ roughly indicates the moment of the (first)
collision. $M_{\rm tot}$ represents the total mass of the initial binary.
In Section 2.2, we describe how we identify the SPH particles that belong to the merger remnant(s). Figure 5 presents the results of precisely calculating the mass loss fraction using the iterative Bernoulli method. It is evident that the fates of most particles - whether they remain part of the merger remnant or are lost - are determined quickly after the collision. This occurs more rapidly in collisions during gentle encounters compared to those in deep encounters for head-on collisions, as the higher impact velocity in deep encounters results in a more violent interaction among fluid elements, leading to a slower mass loss process. For grazing collisions in the gentle encounter (Run 4), the non-monotonic behavior of the mass loss arises from a sequence of successive collisions. After the final collision (at $t\gtrsim 40\,t_{\rm dyn}$; see also Figure 8), the mass loss fraction gradually settles down to a stable value.
@@
For head-on collisions in both deep and gentle encounters, the mass loss fraction stabilizes at approximately 5%. For grazing collisions, the mass loss fraction is more sensitive to the impact parameter, with less mass loss for larger impact parameters.
Figure 6: The density profile of the merger remnant (left column) and the fraction of the total mass enclosed within a given radius relative to the total mass of the merger remnant (right column) are shown for Run 1a (upper row) and Run 3 (lower row). The solid lines represent the profiles at different times after the collision, while the dashed line corresponds to a $2M_{\odot}$ $\gamma=5/3$ polytrope with radius determined by the mass-radius relation $R_{\star}\propto M_{\star}^{0.8}$, which also approximates the structure of a MS star. The insets in the right panels highlight the mass distribution at the outskirts of the remnant.
In addition to the mass loss, we are also interested in the structural properties of the merger remnant. Clearly, not all particles settle down immediately after the collision, and some asymmetry exists within the merger remnant during the short post-collision phase. For now, we ignore the effects of this asymmetry. Starting from the remnant’s center-of-mass, we calculate the density $\rho_{\rm rem}$ at different radius $r_{\rm rem}$ by dividing the mass and volume of spherical shells. We also compute the fraction of the total mass enclosed within a given radius $M_{\rm rem}(<r_{\rm rem})$ relative to the total mass of the merger remnant, as determined using the iterative Bernoulli method. Figure 6 illustrates the structure of the merger remnant following the head-on collision (Runs 1 & 3).
@@
Compared to a typical $2M_{\odot}$ MS star (modeled here as a $\gamma=5/3$ polytrope, with its radius determined by the mass-radius relation $R_{\star}\propto M_{\star}^{0.8}$; represented by the dashed line in Figure 6), the merger remnant exhibits a more concentrated core and an extremely extended envelope. Over 5% of the total mass is distributed in the envelope at $r_{\rm rem}>5R_{\odot}$.
We can also calculate the total energy of the remnant based on the SPH simulation output. Taking the head-on collision in the deep encounter (Run 1a) as an example, we compute the kinetic energy $E_{k,\rm rem}$ relative to the center-of-mass, gravitational potential energy $E_{p,\rm rem}$, and internal energy $U_{\rm rem}$ for all SPH particles belonging to the remnant. These components were then used to determine the total energy $E_{tot,\rm rem}$ of the remnant. The results are presented in Figure 7.
Figure 7: Energy of the merger remnant in Run 1a as a function of time. The different solid lines represent the kinetic energy $E_{k,\rm rem}$ (relative to the remnant center-of-mass), gravitational potential energy $E_{p,\rm rem}$, internal energy $U_{\rm rem}$, and total energy $E_{\rm tot,\rm rem}$.
@@
For the grazing collisions (Runs 2 & 4), the larger impact parameters typically prevent the two stars from merging during the first collision. In the case of the deep encounter (Run 2), the tidal forces from the SMBH have already disrupted the binary energy-wise. After the first collision, the two stars do not collide again but instead undergo a process resembling the standard binary tidal breakup: one star becomes bound to the SMBH’s orbit, while the other is ejected. However, both stars have undergone significant perturbations, leaving their structures vastly different from those of typical MS stars.
@@
Figure 9: Same as Figure 6, but for Runs 2 and 4. Note that the “bump” in the blue line in the upper left panel represents the other remnant.
We also analyze the property of the merger remnant following a grazing collision (Runs 2 & 4), with the results shown in Figure 9. Unlike the cases of head-on collisions, for the deep encounter, we select the remnant that is ultimately bound to the SMBH and compared it to a $1M_{\odot}$ $\gamma=5/3$ polytrope.
@@ 3.2 Mass Loss Debris and Fallback
In Paper I, we suggested the mass loss from stellar collisions induced by binary-SMBH encounters will accrete onto the SMBH, generating accretion luminosity or flares even without a stellar TDE. As shown in Figures 1-5, it is indeed the case that, since the collision velocity is typically comparable to the stellar surface escape velocity (see Figure 9 of Paper I), collisions generally result in a mass loss of approximately 5%. This is consistent with the results of stellar collisions in the absence of an SMBH (e.g., Benz & Hills, 1987; Lai et al., 1993; Freitag & Benz, 2005). A key difference between this mass-loss debris and the typical TDE debris is that the former is not produced by tidal forces, and thus does not remain in the same plane or have the very small launch angles characteristic of the TDE debris (which results in an extremely narrow stream). Instead, the mass-loss debris generated by stellar collisions typically has much larger launch angles, and therefore, in terms of morphology, these debris are more like a “cloud” than a “stream”. Similar to TDEs, because our binary is initially on a parabolic orbit relative to the SMBH, only about half of the mass-loss debris will fall back onto the SMBH.
Figure 10: Probability density function (PDF) of the orbital energy of the mass-loss debris relative to the SMBH after collisions. The results are scaled by $\varepsilon_{\rm col}$, as defined in Eq. (16).
@@
From Figure 10, we see that for collisions that result in a single remnant (Runs 1a, 3, and 4), the debris energy distribution exhibits a bimodal structure with a relatively long tail. The bimodal feature arises because the orbital energy of the merger remnant relative to the SMBH is nearly zero, so most particles with nearly zero orbital energy remain bound to the merger remnant. The long tail indicates that many debris particles have relatively high launch velocities relative to the merger remnant. This is due to the shock waves and fluid forces present during the collision, allowing some debris to achieve higher energies relative to the merger remnant. For a grazing collision in a deep encounter that results in two perturbed stars (Run 2; upper-right panel in Figure 10), the energy distribution exhibits a triple-peaked structure in addition to the long tail. This can be interpreted in a similar manner as the bimodal cases, where debris particles around the dips are bound to the corresponding stellar remnants. In this case, however, the remnants themselves have non-zero orbital energies—approximately symmetric and opposite in sign—corresponding to the two dips near $\varepsilon_{\rm col}\sim\pm 2$. We also see that in our representative examples, the collisions in deep encounter (Run 1a , 2) results in a wider energy distribution compared to the collision in the gentle encounter (Run 3 , 4).
@@
Figure 11: Fallback accretion rate, $\dot{M}(t)$, onto the SMBH. The different colored lines in the upper left panel represent the results for different particle numbers ($N=5\times 10^{5}$ and $10^{5}$ per star for Run 1a and 1b). To convert to physical units, we have set $M_{\rm BH}=10^{6}M_{\odot}$, $M_{\star}=1M_{\odot}$ and $R_{\star}=1R_{\odot}$. The black dashed line shows the standard mass fallback rate for a $1M_{\odot}$ star undergoing a TDE, while the red dashed line represents a power-law curve that matches the late-time fallback rate. The dotted horizontal line indicates the Eddington limit for a $10^{6}M_{\odot}$ black hole, assuming an accretion efficiency of $\epsilon=0.1$. The blue dotted vertical line in the upper right panel marks the moment when the merger remnant returns to the pericenter of its orbit around the SMBH. Here, $t=0$ corresponds to the moment of stellar collision.
@@
The fallback rate curves of the mass-loss debris differ significantly from the standard TDE curve. These curves exhibit a slow rise over an extended period, remain relatively steady, and then follow a power-law decline similar to that of a TDE. For the grazing collision in a deep encounter, as previously discussed, debris with specific orbital binding energy similar to that of the bound remnant tends to remain gravitationally bound and effectively becomes part of the remnant. This is reflected in the dip appearing around $\sim 250$ days in the curve (Figure 11, the upper right panel), which coincides with the time when the remnant returns to the pericenter of its orbit around the SMBH. The relationship between the curve shape and $\beta_{b}$ is not very clear. The peak fallback rate is much lower than that of a TDE, as the total fallback mass is only a few percent of the fallback mass in a TDE for a star of the same mass. However, the minimum fallback time for the debris is significantly shorter compared to the TDE case, particularly for deep encounters (the upper row). In gentle encounters, this effect is less pronounced, as the collision occurs long after the pericenter passage. Despite comparable total mass loss, there is a significant difference in the magnitude of the peak fallback rates. In the deep encounter (Run 1 , 2), the collision occurs very close to the pericenter, causing the debris to fall back onto the SMBH more quickly. This results in a substantially higher peak rate, which remains above the Eddington limit for several hundred days. In contrast, for the gentle encounter (Run 3 , 4), the fallback occurs much more slowly, and the curve consistently stays below the Eddington limit.
@@
We can estimate the approximate size of the final disk by calculating the pericenter distances of the mass-loss SPH particles, based on their angular momenta (relative to the SMBH) after the head-on collision s (Runs 1a, 3). The results are shown in Figure 12.
@@
4 Results: MESA stars
The $\gamma=5/3$ polytrope is considered in section 3 underestimates the central density concentration of solar-type MS stars. As the hydrodynamics of stellar collisions is highly sensitive to the initial stellar mass distribution (e.g., Lai et al., 1993), we perform a comparative set of simulations using a more realistic $M=M_{\odot}$ MS star model generated by the stellar evolution code MESA. For simplicity and to maintain a clear comparison with the polytropic case presented in section 3, we focus on head-on and grazing collisions in the deep encounter regime.
The overall collisional dynamics share many qualitative similarities with the polytropic results. The most striking difference, however, lies in the outcome of the grazing encounter. In the polytropic case, the two stars experience a single pericenter passage and continue on unbound hyperbolic trajectories as two separate, highly perturbed stars. For the MESA stars, we find the different outcome is driven by the initial collision parameters. A comparison of Table 1 shows that the relative velocity for the MESA star encounter (Run 6) is lower than for the polytropic case (Run 2). This lower kinetic energy allows the stars to become gravitationally bound. Consequently, they are unable to escape each other and are quickly drawn back for a second, decisive encounter that leads to a full merger into a single remnant. This process of tidal capture followed by a merger is analogous to the behavior seen in the gentle grazing encounter of polytropic stars depicted in Figure 8.
We now analyze three key diagnostics as in Section 3: the mass loss fraction, the post-merger remnant structure, and the fallback accretion rate of mass-loss debris.
Figure 13: Same as Figure 5, but for Runs 5 and 6.
Figure 13 presents the mass loss fractions calculated using the iterative Bernoulli method for Runs 5 and 6. Similar to the polytropic cases, the mass loss fractions for MESA stars also stabilize shortly after the collision. The stellar structures are different for the two collision types. The head-on collision (Run 5) ejects more mass than its polytropic equivalent. This is because the highly concentrated cores of the MESA stars collide directly, leading to stronger shocks and a more energetic explosion that unbinds a larger fraction of the stellar material. Conversely, the grazing collision (Run 6) ejects less mass. Here, the interaction is confined to the low-density outer envelopes of the MESA stars, while the dense, tightly bound cores miss each other. This less disruptive encounter unbinds a smaller amount of material.
Figure 14: Same as Figure 6, but for Runs 5 and 6. The dashed line represents the initial $1\,M_{\odot}$ MAMS profile obtained from MESA. The bump in the lower-left panel indicates the presence of the secondary star’s core during the final inspiral phase, just before the full merger.
The internal structure of the final merger remnants is detailed in Figure 14. A key feature, common to both head-on and grazing collisions, is that the remnant’s central density is lower than that of the initial progenitor star, even though its total mass is nearly doubled. The immense kinetic energy dissipated during the collision causes the core to expand, resulting in a “puffed-up” thermally supported remnant. This core is surrounded by an extended envelope, qualitatively matching the structure seen in the polytropic mergers. Quantitatively, the remnant radii are slightly more compact than their polytropic counterparts, though they are still in a phase of slow, post-merger expansion driven by thermal relaxation.
Figure 15: Same as Figure 11, but for Runs 5 and 6.
Finally, Figure 15 illustrates the mass fallback rates following the collisions. The magnitude of the fallback rate is correlated with the total mass loss. The head-on collision (Run 5), with its higher mass loss, exhibits a higher peak fallback rate. The grazing collision (Run 6) shows a lower rate. The temporal evolution also generally follows a power-law decay. A significant feature is the smooth fallback curve for the grazing merger (Run 6) - the grazing collision in this case leads to a single remnant. Hence, the fallback rate curve does not exhibit a dip as seen in the polytropic grazing case with two remnants (cf. Figure 11, upper right panel).
@@ 5.1 Summary of Key Results
In this paper, we have conducted a detailed study of the stellar collisions induced by the close encounter between a stellar binary and a SMBH. The binary, with a total mass $m_{12}$, an initial circular orbit, and a semi-major axis $a_{b}$, approaches the SMBH (mass $M_{\rm BH}$) on a parabolic orbit, with a pericenter distance $r_{p}$ from the SMBH. Using SPH simulations, we characterize key properties of stellar collisions and their potential observational consequences. We consider two type of stellar models: a $\gamma=5/3$ polytropic model (representing low-mass MS stars with mass $\lesssim 0.4M_{\odot}$) and a solar-like MS star obtained from MESA.
@@
    Mass loss and remnant(s). Since the collision velocity is comparable to the stellar escape velocity (see Table 1 and Figure 9 in Paper I), some mass loss is expected during the collision. Head-on collisions generally produce a single merger remnant, with a few percent mass loss (Figs. 5, 13) In contrast, grazing collisions exhibit more diverse outcomes. For gentle encounters, while the first collision may not immediately result in a single merger remnant, the binary remains gravitationally bound; the perturbed stars undergo multiple subsequent collisions (Figure 8), ultimately forming a single remnant with minimal mass loss ($\lesssim 1\%$). However, in deep encounters, where the binary system is already disrupted, if the first collision fails to produce a merger, two remnants can form, as seen in our Run 2, and the mass loss fraction is relatively higher ($\sim 4\%$). Note that the outcome of grazing collisions in deep encounters is highly sensitive to the collision parameters and stellar structure. While polytropic models result in two separate remnants, simulations with solar-type MESA stars yield a lower relative velocity and a tidal capture, resulting in a single merger remnant. This suggests that for solar-type stars, mergers are a more common outcome than for low-mass MS stars.
@@
    Density structure of merger remnant(s). The remnant star formed after collision has a significant different structure compared to the original MS star. The merger remnant is characterized by a highly extended envelope. Its core structure depends on the progenitor model: for polytropic stars, the remnant has a much higher central density, whereas for $1M_{\odot}$ MESA stars, the dissipated kinetic energy causes the core to expand, resulting in a lower central density than the initial star (Figs. 6, 9, 14). The extended envelope makes the merger remnant more susceptible to partial tidal disruption when it returns to the pericenter around the SMBH. If the two stars do not merge but instead undergo significant dynamical perturbations, the outcomes differ substantially (upper row of Figure 9). These perturbed stars not only develop an extended envelope but also experience a substantial reduction in the central density. Analogous to the Hills mechanism, one of the stars often remains bound to the SMBH. In this case, this perturbed star is also significantly more susceptible to tidal disruption.
@@
    Orbital property of mass-loss debris and mass fallback rate. Nearly half of the mass-loss debris produced during the collision is expected to fall onto the SMBH, undergoing a process analogous to a TDE. From the simulation snapshots (Figs. 1-4), we observe that the “debris cloud” generated by the collision exhibits a morphology distinctly different from the “debris stream” typical of TDEs. Furthermore, an analysis of the binding energy distribution of the debris (Figs. 1013) reveals significant deviations from the top-hat distribution observed in stellar TDEs. Specifically, the energy distribution is bimodal for collisions that produce a single remnant, and becomes triple-peaked for the cases that result in two separate remnants, with the dips corresponding to debris gravitationally bound to the remnant(s). Since the collision does not necessarily occur at the pericenter, unlike in the standard TDE, the resulting mass fallback accretion rate also differs (Figs. 1115). Although the total mass loss represents only a small fraction of the stellar mass and thus results in a relatively low magnitude of fallback accretion, the shape of the fallback rate curve is notably distinct from that of TDEs. Specifically, the fallback curve shows notable features depending on the collision’s outcome. Generally, it remains approximately constant for an extended period ($\gtrsim 100$ days) before starting to decline (for $M_{\rm BH}=10^{6}M_{\odot}$ and solar-type stars). For the two-remnant case arising from a deep grazing collision of polytropes, the fallback rate exhibits a distinct dip, caused by the bound remnant returning to its pericenter and re-capturing some of the debris (Figure 11). In contrast, for the MESA case in this work, which our grazing-encounter yields a single remnant, the fallback curve is smooth and lacks such a feature (Figure 15). Additionally, the angular momentum distribution of the mass-loss debris (represented by the debris’ pericenter distance $r_{p}$, see Figure 12) suggests that the resulting accretion disk extends to the original pericenter distance.